OpenStudy (anonymous):

Rusty has 4 racing video games, 2 trivia video games, and 5 action video games. A friend borrows one of the racing video games. Rusty decides to play one of his remaining video games. If he randomly picks one to play, what is the probability that he picks a racing game?

OpenStudy (anonymous):

after friend had borrowed the no of games remained are 3 racing, 2 trivia, 5action so total no.=10 you have total no. out of what you want something find probability

OpenStudy (anonymous):

No of ways in which he can pick a racing game=3 No of ways in which he can pick *any* game=3+2+5=10 \[\Large\text{Probability of picking a racing game} =\frac{3}{10}\]

OpenStudy (anonymous):

yo!!